OpenStudy (anonymous):

Solve.3 over x minus 1 + 2 over x = 4 A. {0.25, 2} B. {1/2, 2} C. {1/2, 1} D. {.25, 1}

OpenStudy (anonymous):

3/x-1+2/x=4

OpenStudy (lgbasallote):

if 1 is also under 3 together with x, then multiply both sides by the LCM first which is x(x-1) so 3x + 2(x-1) = 4(x(x-1)) 3x + 2x -2 = 4x^2 - 4x -4x^2 + 4x + 3x +2x = 2 -4x^2 + 9x = 2 x(-2x + 9) = 2 x = -2; x = 3.5

OpenStudy (anonymous):

oh okay i was confused about the lcm thank you